# Learning from Negative Samples in Generative Biomedical Entity Linking

### v1

@@ Abstract
Generative models have become widely used in biomedical entity linking (BioEL) due to their excellent performance and efficient memory usage. However, these models are usually trained only with positive samples—entities that match the input mention’s identifier—and do not explicitly learn from hard negative samples, which are entities that look similar but have different meanings. To address this limitation, we introduce ANGEL (Learning from Negative Samples in Generative Biomedical Entity Linking), the first framework that trains generative BioEL models using negative samples. Specifically, a generative model is initially trained to generate positive samples from the knowledge base for given input entities. Subsequently, both correct and incorrect outputs are gathered from the model’s top-k predictions. The model is then updated to prioritize the correct predictions through direct preference optimization. Our models fine-tuned with ANGEL outperform the previous best baseline models by up to an average top-1 accuracy of 1.4% on five benchmarks. When incorporating our framework into pre-training, the performance improvement further increases to 1.7%, demonstrating its effectiveness in both the pre-training and fine-tuning stages. Our code is available at https://github.com/dmis-lab/ANGEL.
Learning from Negative Samples in Generative Biomedical Entity Linking
@@ 1 Introduction
Biomedical entity linking (BioEL) involves aligning entity mentions in text with standardized concepts from biomedical knowledge bases (KB) such as UMLS Bodenreider (2004) or MeSH Lipscomb (2000). BioEL encounters significant challenges due to the diverse and ambiguous nature of biomedical terminology, including synonyms, abbreviations, and terms that look similar but have different meanings. For instance, ‘ADHD’ (CUI:C1263846, where CUI stands for Concept Unique ID) has synonyms such as hyperkinetic disorder and attention deficit hyperactivity disorder. Additionally, ‘ADA’ can be mapped to either adenosine deaminase (CUI:C1412179) or American Diabetes Association (CUI:C1705019) depending on the context in which the entity appears.
Recent studies have focused on addressing these challenges, broadly categorized into two approaches: similarity-based and generative BioEL. Similarity-based models Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Bhowmik et al. (2021); Agarwal et al. (2022) encode input mentions and entities from KBs into the same vector space using embedding models. They then calculate similarity scores to identify the most similar entities for each input entity. Although these approaches have achieved remarkable improvements, they require significant space to index and load embedding vectors for all candidate entities De Cao et al. (2020). Furthermore, representing both the input and candidate entities as single vectors using a bi-encoder can limit the quality of their representations, making it difficult to handle challenging cases.
On the other hand, generative models De Cao et al. (2020); Yuan et al. (2022b), built upon an encoder-decoder structure Lewis et al. (2020); Raffel et al. (2020), directly generate the most likely entity name from the KB for the input entity. The output space is dynamically controlled through a constrained decoding strategy, ensuring that only entities from the target KB are generated. Generative models offer several advantages over similarity-based models, including greater memory efficiency and higher performance. They eliminate the need to index large external embedding vectors, and their auto-regressive formulation effectively cross-encodes the input document and candidate entities.
@@
To harness the benefits of generative approaches while overcoming their limitation of not using negative samples, we introduce a novel training framework, ANGEL. Our framework operates in two stages: positive-only training and negative-aware training. In the first stage, a generative model is trained to generate biomedical terms from the KB that share the same identifier as the given input entity. In the second stage, we gather both correct and incorrect outputs from the model’s top-k predictions. The model is then updated to prioritize the correct predictions using the direct preference optimization (DPO) algorithm Rafailov et al. (2024). Models trained on our ANGEL framework significantly outperform the previous best similarity-based and generative BioEL models, achieving an average accuracy improvement of 1.7% across five datasets. Our contributions are as follows:
@@ 2.2 Generative Entity Linking
Generative models have become a powerful method for entity linking by overcoming the limitations of similarity-based models. The GENRE framework De Cao et al. (2020) was the first to demonstrate this approach. To enhance precision and reduce memory usage, GENRE introduced a constrained decoding method Hokamp and Liu (2017) using a prefix tree (trie), which restricts the output space to valid entity names. This technique also facilitates easy updates to the set of entities, making the system highly adaptable to changes in the KB. In the biomedical field, notable examples of generative models include GenBioEL Yuan et al. (2022b) and BioBART Yuan et al. (2022a). GenBioEL, in particular, is the first model to apply a generative model BART Lewis et al. (2020) to BioEL, after pre-training it using UMLS. Additionally, several hybrid approaches, known as retrieve-and-generate methods, have been proposed Xu et al. (2023); Lin et al. (2024). In these methods, a similarity-based model first retrieves the top-k candidates, which are then reranked using a generative model. Although generative approaches have shown high performance, their training has typically been limited to positive samples, as discussed in the introduction section. In this study, we introduce the use of negative samples during training and demonstrate that this approach can significantly enhance the performance of generative models.
@@ 3.1 Task Formulation
Let be a human-labeled dataset, where represents an input text and is the gold identifier defined in a KB denoted by . Each contains a target entity mention along with its surrounding contextual information and , which represents the tokens before and after the entity mention , respectively. For simplicity, we will omit the subscript . Our goal is to map each mention to its corresponding identifier from the set of entity names as follows:
where is a mapping function that converts entities to their identifiers, and represents the model parameters.
Previous generative BioEL approaches train the model to generate a synonym for the given mention in an autoregressive manner as follows:
where is the set of entity names (i.e., synonyms) corresponding to the identifier , and is the number of tokens of the synonym and indicates the -th token of the synonym. The prefix prompt to the decoder, represented as ‘[BOS] is’, is designed to make the decoder’s output resemble a natural language sentence, which helps to minimize discrepancies between language modeling and fine-tuning on the BioEL task. The target mention in the input is surrounded by the special tokens, [ST] and [ET], as follows:
where the special tokens [BOS] and [EOS] represent the ‘Begin Of Sentence’ and ‘End Of Sentence’, respectively.
As shown in Equation 2, existing models are trained only to output synonyms corresponding to the given mention (i.e., positive samples), without learning from negative samples. In contrast, we introduce a new method called negative-aware training, which allows the model to learn by comparing both positive and negative samples, enhancing the model’s generalizability. We will describe our framework in detail in the following sections.
@@ 3.2 ANGEL Framework
Our framework consists of two main stages: positive-only training, which warms up the model on target datasets, and negative-aware training, which continuously improves the model by learning from negative samples (see Figure 2).
@@
In this initial stage, the goal is to learn the morphological similarities among synonyms. To achieve this, we train the model to generate synonyms (i.e., positive samples) that are predefined in the KB, similar to traditional generative methods (see Equation 2). Our preliminary study indicated that using all synonyms, , was not effective. Also, relying solely on the top-1 synonym, as done in a previous study Yuan et al. (2022b), may limit generalizability. Therefore, we select several of the most similar synonyms to the mention based on their TF-IDF similarity, which is calculated as follows:
where returns similarity scores. We use the top-k subset as training instances for each mention.
@@
Although surface similarities are a useful feature for BioEL, over-relying on them can limit the model’s generalization ability. To address this issue, we update the model using negative-aware training. First, we obtain the top-k predictions of the model for mentions in the training dataset. We then automatically construct a training dataset consisting of triplets: a mention with context (if it exists), a correct prediction, and an incorrect prediction. Among all pairs of correct and incorrect predictions, we select only those pairs where the incorrect prediction’s rank is higher than the correct prediction’s rank. Particularly, when the highest ranked entity is the correct one, we pair this entity with the highest ranked incorrect entity to preserve the model’s prior learning. Finally, using this dataset , we then optimize the model with the DPO algorithm. This maximizes the likelihood of generating the correct prediction, , over the incorrect prediction, , defined as follows:
where is the generative model to be trained, is the generative model that was trained in the previous stage using positive-only training, is a sigmoid function, and is a hyperparameter.
@@
Our framework is versatile, supporting not only fine-tuning with labeled datasets but also pre-training with the KB itself. Initially, we conduct positive-only training using synonym lists defined in UMLS, the most extensive KB in the biomedical field. We generate its surrounding contextual information automatically for each entity, using clause templates or definitions, as outlined in GenBioEL Yuan et al. (2022b).111Detailed descriptions of data generation during the pre-training stage are provided in the appendix. We then use the TF-IDF similarity to identify the top synonym and set it as the target (Equation 3). For negative-aware training, using the model’s top-k predictions is impractical due to UMLS’s vastness, which includes over 3 million entities. Instead, we use entities with the highest TF-IDF similarity but different identifiers from the input mentions as negative samples. This approach significantly reduces computation, enabling effective training across the entire UMLS.
@@
| Similarity-based BioEL | ||||||
| BioSYN Sung et al. (2020) | 91.1 | - | 71.3 | 82.6 | - | - |
| SapBERT Liu et al. (2021) | 92.3 | - | 75.1 | 89.0 | - | - |
| ResCNN Lai et al. (2021) | 92.4 | - | 80.1 | - | 55.0 | - |
| KRISSBERT Zhang et al. (2022) | 91.3 | - | - | - | 72.2 | - |
@@ 4.1 Datasets
We utilized five popular BioEL benchmark datasets: NCBI-disease Doğan et al. (2014), BC5CDR Li et al. (2016b), COMETA Basaldella et al. (2020), AskAPatient Limsopatham and Collier (2016b), and MedMentions Mohan and Li (2019), with the ST21pv subset used for MedMentions. Due to the lack of a test set in the AskAPatient dataset, we adhere to the 10-fold evaluation protocol outlined by Limsopatham and Collier (2016b). Also, this dataset does not include context for the mentions. In the following tables, NCBI-disease, AskAPatient, and MedMentions are denoted as NCBI, AAP, and MM-ST21pv, respectively. Please refer to the appendix for detailed dataset descriptions and statistics.
@@ 4.2 Baseline Models
We use top-performing similarity-based models Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Zhang et al. (2022) and generative models Lewis et al. (2020); Yuan et al. (2022a, b) as our baselines. To ensure a fair comparison with our model under identical experimental conditions, we replicate the following generative models and then apply our ANGEL framework to them: (1) BART-large Lewis et al. (2020) is an encoder-decoder language model pre-trained on a general-domain corpus. (2) BioBART-large Yuan et al. (2022a) is the BART-large model continuously pre-trained on a biomedical-domain corpus. (3) GenBioEL Yuan et al. (2022b) is initialized with the weights of the BART-large model and then pre-trained specifically for BioEL using UMLS.
In BioEL, several studies utilize retrieve-and-generate methods Xu et al. (2023); Lin et al. (2024). This involves a similarity-based model retrieving the top-k candidates from the KB, followed by a generative model reranking these candidates. We exclude these methods from our experiments to focus on introducing a novel training method that uses negative samples for generative entity linking and demonstrating its effectiveness in a single generative model. Future research could apply our methodology to enhance the reranking process of generative models.
@@ 4.3 Implementation Details
Our framework is applied to each of these models during fine-tuning, referred to as ANGEL, and during both pre-training and fine-tuning, referred to as ANGEL. For pre-training, we utilized the 2020AA version of the UMLS database,222https://www.nlm.nih.gov/research/umls/licensedcontent/umlsarchives04.html which comprises 3.09M entities, of which 199K concepts contain definitions. In the pre-training phase, the model was trained for 5 epochs, with checkpoints created every 500 steps. We selected the best checkpoints based on the validation sets. During fine-tuning on MM-ST21pv, we also used the 2020AA version of UMLS because the 2017AA version was not directly accessible. Please note that the reported scores of baseline models were measured based on the 2017AA version of UMLS. For determining the best hyperparameters in positive-only training, we searched for the optimal learning rate within the range of 1e-5 to 3e-7 and adjusted the batch size between 8 and 16, following the approach of Yuan et al. (2022b). For the negative-aware training, we searched for the optimal learning rate from 2e-5 to 1e-6 and experimented with batch sizes ranging from 16 to 64. In pre-processing, following Yuan et al. (2022b), we expanded abbreviations using AB3P Sohn et al. (2008), lowercase texts, mark mention boundaries with special tokens [ST] and [ET], and discard mentions that overlap or are missing from the target KB. The best hyperparameter configurations are detailed in the Appendix. We used the source codes provided by Yuan et al. (2022b)333https://github.com/Yuanhy1997/GenBioEL and alignment handbook Tunstall et al. (2023)444https://github.com/huggingface/alignment-handbook. During pre-training stage, we trained our model using eight 80G A100 GPUs for 12 hours. During fine-tuning stage, we used a single A100 GPU.
@@ 4.4 Results
Consistent with previous studies Sung et al. (2020); Liu et al. (2021), we use accuracy at top-1 (Acc@1) as our evaluation metric. This metric measures the percentage of mentions where the model correctly ranks the gold standard identifier as the top choice. Table 1 demonstrates that our framework consistently improves the performance of generative models. Specifically, our fine-tuning method (i.e., ANGEL) improves the Acc@1 scores of BART, BioBART, and GenBioEL by 0.9%, 1.4%, and 1.2%, respectively. When pre-training is also applied (i.e., ANGEL) to GenBioEL, the improvement increases to 1.7%, underscoring the effectiveness of both pre-training and fine-tuning in ANGEL. Our best model (i.e., GenBioEL with ANGEL) outperforms all baseline models, whether they are similarity-based or generative BioEL models.

### v2

# Learning from Negative Samples in Biomedical Generative Entity Linking

@@ Abstract
Generative models have become widely used in biomedical entity linking (BioEL) due to their excellent performance and efficient memory usage. However, these models are usually trained only with positive samples, i.e., entities that match the input mention’s identifier, and do not explicitly learn from hard negative samples, which are entities that look similar but have different meanings. To address this limitation, we introduce ANGEL (Learning from Negative Samples in Biomedical Generative Entity Linking), the first framework that trains generative BioEL models using negative samples. Specifically, a generative model is initially trained to generate positive entity names from the knowledge base for given input entities. Subsequently, both correct and incorrect outputs are gathered from the model’s top-k predictions. Finally, the model is updated to prioritize the correct predictions through preference optimization. Our models outperform the previous best baseline models by up to an average top-1 accuracy of 1.4% on five benchmarks. When incorporating our framework into pre-training, the performance improvement increases further to 1.7%, demonstrating its effectiveness in both the pre-training and fine-tuning stages. The code and model weights are available at https://github.com/dmis-lab/ANGEL.
Learning from Negative Samples in Biomedical Generative Entity Linking
@@ 1 Introduction
Biomedical entity linking (BioEL) involves aligning entity mentions in text with standardized concepts from biomedical knowledge bases (KB) such as UMLS Bodenreider (2004) or MeSH Lipscomb (2000).111UMLS and MeSH are short for the Unified Medical Language System and Medical Subject Headings, respectively. BioEL encounters significant challenges due to the diverse and ambiguous nature of biomedical terminology, including synonyms, abbreviations, and terms that look similar but have different meanings. For instance, ‘ADHD’ (CUI:C1263846, where CUI stands for Concept Unique ID) has synonyms such as hyperkinetic disorder and attention deficit hyperactivity disorder. Additionally, ‘ADA’ can be mapped to either adenosine deaminase (CUI:C1412179) or American Diabetes Association (CUI:C1705019) depending on the context in which the entity appears.
Recent studies have focused on addressing these challenges, broadly categorized into two approaches: similarity-based and generative BioEL. Similarity-based models Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Bhowmik et al. (2021) encode input mentions and entities from KBs into the same vector space using embedding models. They then calculate similarity scores to identify the most similar entities for each input entity. Although these approaches have achieved remarkable improvements, they require significant space to index and load embedding vectors for all candidate entities De Cao et al. (2020). Furthermore, representing both the input and candidate entities as single vectors using a bi-encoder can limit the quality of their representations, making it difficult to handle challenging cases.
On the other hand, generative models De Cao et al. (2020); Yuan et al. (2022a, b), built upon an encoder-decoder structure Lewis et al. (2020); Raffel et al. (2020), directly generate the most likely entity name from the KB for the input entity. The output space is dynamically controlled through a constrained decoding strategy, ensuring that only entities from the target KB are generated. Generative models offer several advantages over similarity-based models, including greater memory efficiency and higher performance. They eliminate the need to index large external embedding vectors, and their auto-regressive formulation effectively cross-encodes the input document and candidate entities.
@@
To harness the benefits of generative approaches while overcoming their limitation of not using negative samples, we introduce a novel training framework, ANGEL. Our framework operates in two stages: positive-only training and negative-aware training (see Figure 2). In the first stage, a generative model is trained to generate biomedical terms from the KB that share the same identifier as the given input entity. In the second stage, we gather both correct and incorrect outputs from the model’s top-k predictions. The model is then updated to prioritize the correct predictions using a preference optimization algorithm Burges (2010); Rafailov et al. (2024). Models trained on our ANGEL framework significantly outperform the previous best similarity-based and generative BioEL models, achieving an average accuracy improvement of 1.7% across five datasets. Our contributions are as follows:
@@ 2.2 Generative Entity Linking
Generative models have become a powerful method for entity linking by overcoming the limitations of similarity-based models. The GENRE framework De Cao et al. (2020) was the first to demonstrate this approach. To enhance precision and reduce memory usage, GENRE introduced a constrained decoding method Hokamp and Liu (2017) using a prefix tree (trie), which restricts the output space to valid entity names. This technique also facilitates easy updates to the set of entities, making the system highly adaptable to changes in the KB. In the biomedical field, notable examples of generative models include GenBioEL Yuan et al. (2022b) and BioBART Yuan et al. (2022a). GenBioEL, in particular, is the first model to apply a generative model BART Lewis et al. (2020) to BioEL, after pre-training it using UMLS. Additionally, several hybrid approaches, known as retrieve-and-generate methods, have been proposed Xu et al. (2023); Lin et al. (2024). In these methods, a similarity-based model first retrieves the top-k candidates, which are then reranked using a generative model. Although generative approaches have shown high performance, their training has typically been limited to positive samples, as discussed in the introduction section. This absence of explicit negative sample learning often leads to confusion when entities share very similar surface forms but represent different concepts. In this study, we introduce the use of negative samples during training and demonstrate that this approach can significantly enhance the performance of generative models.
@@ 3.1 Task Formulation
Let be a human-labeled dataset, where represents an input text and is the gold identifier defined in a KB denoted by . Each contains a target entity mention along with its surrounding contextual information and , which represents the tokens before and after the entity mention , respectively. For simplicity, we will omit the subscript . Our goal is to map each mention to its corresponding gold identifier from the set of entity names . To achieve this, we define the model’s prediction as follows:
where is an entity name defined in the KB, is a function that aligns entities to their identifiers, and represents the model parameters.
A single gold identifier may have multiple associated entity names that refer to the same concept; we refer to these as synonyms. Previous generative BioEL approaches train the model to generate a textual synonym , where denotes the set of entity names associated with the identifier , in an autoregressive manner as follows:
where is the number of tokens of the synonym , indicates the -th token of the synonym, and is the prompt. In an encoder-decoder model structure Lewis et al. (2020), the input to the encoder is formatted as follows:
where the special tokens [ST] and [ET] surround the target mention, and the special tokens [BOS] and [EOS] represent the ‘Begin Of Sentence’ and ‘End Of Sentence,’ respectively. The prefix prompt to the decoder, represented as ‘ is’, is concatenated with [BOS] and input to the decoder. The prompt is designed to make the decoder’s output resemble a natural language sentence, which helps to minimize discrepancies between language modeling and fine-tuning on the BioEL task.
As shown in Equation 2, existing models are trained solely to predict synonyms for the input mention (i.e., positive samples), without leveraging negative samples. In contrast, we propose a novel approach using negative samples, which we will describe in detail in the following sections.
@@ 3.2 ANGEL Framework
Our framework comprises two main stages: positive-only training, which warms up the model using positive samples to learn morphological similarities among synonyms, and negative-aware training, which progressively refines the model by incorporating negative samples (see Figure 2).
@@
We initialize the model to generate synonyms, similar to previous methods (see Equation 2). For the input mention, we select the most similar synonyms based on their vector similarity, which is calculated as follows:
where returns the TF-IDF similarity between tri-grams of the mention and its synonyms. We use the top-k subset as training instances for each mention.
@@
After obtaining the top-k predictions from the model for each mention in the training set, we construct a dataset of triplets , where denotes the mention along with its context (if available), is the correct (preferred) entity, and is an incorrect (dispreferred) entity. From all possible pairs, we retain only those for which the model ranks the incorrect entity above the correct one , thereby reflecting an incorrect model preference. If the top-ranked prediction is already correct, we pair this correct entity with the highest-ranked incorrect entity to preserve the model’s original preference structure.
We denote the resulting training set by and fine-tune the model using a pairwise preference loss, formulated as follows:
where is a differentiable scoring function (e.g., ), is the sigmoid function, and is a temperature (scaling) hyperparameter. A recent instantiation of this general preference learning framework is Direct Preference Optimization (DPO) Rafailov et al. (2024), where the scoring function is defined as a log-ratio with respect to a reference model as follows:
where is the generative model being trained, is a reference generative model trained in a prior stage using positive-only data. We adopt DPO as a practical instantiation within our framework because it offers a principled and empirically effective way to incorporate prior model behavior through a reference distribution.
@@
Our framework supports not only fine-tuning with labeled datasets but also pre-training with the KB. Specifically, we automatically generate surrounding contextual information for each entity in the KB, using clause templates or definitions, as outlined in GenBioEL Yuan et al. (2022b). When the entity definition corresponding to the identifier is available, a synonym and its definition are integrated into a pre-defined clause template as follows:222Refer to Yuan et al. (2022b) for the full set of templates.
When no definitions are available in the KB, we replace with alternative synonyms as follows:
where and are different synonyms. The input for the decoder is “ (or ) is” and the expected output is another synonym (e.g., ) selected from the remaining synonyms.
The pre-training process, like fine-tuning, is divided into two stages: positive-only training and negative-aware training. However, due to the typically large scale of the KB, efficiency considerations are particularly important. In positive-only training, rather than utilizing all possible synonym combinations within the KB, we identify, for each entity, the most similar synonym based on TF-IDF similarity and designate it as the target synonym. For negative-aware training, instead of selecting negatives from the model’s predictions, negative samples are selected from entities exhibiting the highest TF-IDF similarity to the input mentions but possessing distinct identifiers.
@@
Similarity-based BioEL & Re-ranking | ||||||
BioSYN Sung et al. (2020) | 91.1 | 93.3† | 71.3 | 86.5† | OOM | - |
SapBERT Liu et al. (2021) | 92.3 | 88.6† | 75.1 | 89.0 | 50.3† | 79.1 |
ResCNN Lai et al. (2021) | 92.4 | 94.0† | 80.1 | 77.4† | 55.0 | 79.3 |
KRISSBERT Zhang et al. (2022) | 91.3 | 72.0† | 80.1† | 83.1† | 72.2 | 79.7 |
Prompt-BioEL Xu et al. (2023) | 91.9† | 94.3† | 82.7† | 89.7† | 72.6† | 86.2 |
@@ 4.1 Datasets
We utilized five popular BioEL benchmark datasets: NCBI-disease Doğan et al. (2014), BC5CDR Li et al. (2016b), COMETA Basaldella et al. (2020), AskAPatient Limsopatham and Collier (2016b), and Medmentions Mohan and Li (2019), with the ST21pv subset used for Medmentions. Due to the lack of a test set in the AskAPatient dataset, we adhered to the 10-fold evaluation protocol outlined by Limsopatham and Collier (2016b). Also, AskAPatient dataset does not include context for the mentions. In the following tables, NCBI-disease, AskAPatient, and Medmentions are denoted as NCBI, AAP, and MM-ST21pv, respectively. Refer to Appendix A for detailed descriptions and statistics.
@@ 4.2 Baseline Models
We used top-performing similarity-based models Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Zhang et al. (2022) as our baselines. Notably, Prompt-BioEL Xu et al. (2023) employs a re-ranking-based approach. In the first stage, a similarity-based model, such as SapBERT, retrieves the top-k candidate entities from the knowledge base. In the second stage, these candidates are reranked using a cross-encoder. Although Prompt-BioEL may not be directly comparable, as it incorporates additional modules on top of existing models, we report its performance alongside for reference. Additionally, we include the previously best-performing generative models for comparison. (1) BART-large Lewis et al. (2020) is an encoder-decoder language model pre-trained on a general-domain corpus. (2) BioBART-large Yuan et al. (2022a) is the BART-large model continuously pre-trained on a biomedical-domain corpus. (3) GenBioEL Yuan et al. (2022b) is initialized with the weights of the BART-large model and then pre-trained specifically for BioEL using UMLS. We excluded several models Agarwal et al. (2022); Lin et al. (2024) due to the lack of publicly available code or the difficulty in reproducing their reported performance.
@@ 4.3 Implementation Details
Our framework was applied to each of these models during fine-tuning, referred to as ANGEL, and during both pre-training and fine-tuning, referred to as ANGEL. For pre-training, we utilized the 2020AA version of the UMLS database,333https://www.nlm.nih.gov/research/umls/licensedcontent/umlsarchives04.html which comprises 3.09M entities, of which 199K concepts contain definitions. During pre-training, we saved checkpoints every 500 steps over 5 epochs and selected the best one based on the validation sets. We used top-3 synonyms as positive samples in positive-only training. The other hyperparameter configurations are detailed in Appendix B. In pre-processing, following Yuan et al. (2022b), we expanded abbreviations using AB3P Sohn et al. (2008), lowercase texts, mark mention boundaries with special tokens [ST] and [ET], and discard mentions that overlap or are missing from the target KB. During pre-training, our models were trained using eight 80G A100 GPUs for 12 hours. During fine-tuning, a single A100 GPU was used.
@@ 4.4 Results
Consistent with previous studies Sung et al. (2020); Liu et al. (2021), we used accuracy at top-1 (Acc@1) as our evaluation metric, which quantifies the percentage of mentions where the model correctly ranks the gold standard identifier as the top choice. To assess statistical significance, we employed bootstrapping with the same sample size as the original datasets, repeating the process 100 times, followed by a paired t-test. Table 1 shows that our framework consistently outperformed the performance of generative models ( for all comparisons). Specifically, our fine-tuning method (i.e., ANGEL) improved the Acc@1 scores of BART, BioBART, and GenBioEL by 0.9%, 1.4%, and 1.2%, respectively. When pre-training is also applied (i.e., ANGEL) to GenBioEL, the improvement increases to 1.7%, further highlighting the effectiveness of both pre-training and fine-tuning in ANGEL.
Similarity-based models often exhibit limited robustness, with performance varying significantly across datasets. In contrast, generative models tend to deliver more consistent results, highlighting a key strength. Among the baseline methods, the re-ranking-based model Prompt-BioEL achieves strong performance, substantially outperforming its underlying retriever, SapBERT, though at the cost of increased inference time. Notably, our ANGEL model surpasses Prompt-BioEL across all datasets without relying on any re-ranking component, achieving an average improvement of 0.5%. Given this strong baseline, incorporating a re-ranking component into our model in future work may further enhance performance.
Model | NCBI | BC5CDR | COMETA | AAP | MM-ST21pv | Average |
Models with Negative-aware Training | ||||||
ANGEL (Ours) | 92.8 | 94.5 | 82.8 | 90.2 | 73.3 | 86.7 |
Prediction-based TF-IDF-based | 91.8 | 94.4 | 81.6 | 90.0 | 71.5 | 85.9 |
Pairs All Possible Pairs | 92.9 | 94.0 | 81.9 | 90.0 | 72.0 | 86.2 |
within Top-5 Top-10 Predictions | 92.5 | 94.0 | 82.1 | 89.6 | 72.6 | 86.2 |
Models without Negative-aware Training | ||||||
GenBioEL Yuan et al. (2022b) | 91.0 | 93.1 | 80.9 | 89.3 | 70.7 | 85.0 |